Melissa's mother is three years more than twice as old as Melissa. How old is Melissa if her mother is 33?
Answer by ankor@dixie-net.com(22740)

Melissa's mother is three years more than twice as old as Melissa.
How old is Melissa if her mother is 33?
:
let m = Melissa's age
:
" Melissa's mother is three years more than twice as old as Melissa."
2m + 3 = Mothers age
Mother's age is given as 33
therefore
2m + 3 = 33
2m = 33 - 3
2m = 30
m = 30/2
m = 15 yrs is Melissa's age